Three charges line vertically respectively61, -32, -8. 61 and -32 is separated by 50 while -32 and -8 is separated by71. What is magnitude and direction of the total net  force experience by -8?



1
Expert's answer
2021-01-14T10:37:14-0500
"F=k\frac{q_1q_3}{r_1^2}+k\frac{q_2q_3}{r_2^2}\\
F=9\cdot10^9\frac{(61)(-8)}{(50+71)^2}+9\cdot10^9\frac{(-32)(-8)}{71^2}=1.57\cdot10^8N"

Direction: downwards.